# Diffusion & Adversarial Schrödinger Bridges via Iterative Proportional Markovian Fitting

### v4

@@ Abstract
The Iterative Markovian Fitting (IMF) procedure, which iteratively projects onto the space of Markov processes and the reciprocal class, successfully solves the Schrödinger Bridge (SB) problem. However, an efficient practical implementation requires a heuristic modification—alternating between fitting forward and backward time diffusion at each iteration. This modification is crucial for stabilizing training and achieving reliable results in applications such as unpaired domain translation. Our work reveals a close connection between the modified version of IMF and the Iterative Proportional Fitting (IPF) procedure—a foundational method for the SB problem, also known as Sinkhorn’s algorithm. Specifically, we demonstrate that the heuristic modification of the IMF effectively integrates both IMF and IPF procedures. We refer to this combined approach as the Iterative Proportional Markovian Fitting (IPMF) procedure. Through theoretical and empirical analysis, we establish the convergence of the IPMF procedure under various settings, contributing to developing a unified framework for solving SB problems. Moreover, from a practical standpoint, the IPMF procedure enables a flexible trade-off between image similarity and generation quality, offering a new mechanism for tailoring models to specific tasks.
@@ 1 Introduction
Diffusion Bridge models inspired by the Schrödinger Bridge (SB) theory, which connects stochastic processes with optimal transport, have recently become powerful approaches in biology (tong2024simulation; bunne2023schrodinger), chemistry (somnath2023aligned; igashovretrobridge; kim2024discrete), computer vision (liu20232; shi2023diffusion; ksenofontov2025categorical) and speech processing (chen2023schrodinger). Most of these applications deal with either supervised domain translation, e.g., image super-resolution and inpainting (liu20232) or with unpaired translation, e.g., image style-transfer (shi2023diffusion) or single-cell data analysis (tong2024simulation).
@@
This work specifically focuses on unpaired domain translation (zhu2017unpaired, Fig. 2). In this setting, given two domains represented solely by unpaired samples, the goal is to transform a sample from the input domain into a sample related to it in the target domain. In this context, researchers usually use SB-based algorithms because they enforce two key properties: the optimality property, ensuring similarity between the input and the translated object, and the marginal matching property, ensuring the translation of the input domain to the target domain. The motivation for relying on such specialized methods, rather than general text-to-data models, is further discussed in Appx. B.
@@
Early works (de2021diffusion; vargas2021solving; chen2021likelihood; pavon2021data) on using the SB for unpaired domain translation employed the well-celebrated Iterative Proportional Fitting (IPF) procedure (kullback1968probability), also known as the Sinkhorn algorithm (cuturi2014fast). The IPF procedure is initialized with a simple prior process satisfying the optimality property. It then refines this process iteratively through optimality-preserving transformations until the marginal matching property is achieved. In each iteration, IPF decreases the forward KL-divergence $\text{KL}(q^{*}\|q)$ between the current approximation $q$ and the ground-truth Schrödinger Bridge $q^{*}$. However, in practice, approximation errors may cause IPF to suffer from the “prior forgetting”, where the marginal matching property is achieved but optimality is lost (vargas2024transport; vargas2021solving).
@@
The Iterative Markovian Fitting (IMF) procedure (shi2023diffusion; peluchetti2023diffusion; gushchin2024adversarial) emerged as a promising competitor to IPF. Contrary to IPF, IMF starts from a stochastic process satisfying the marginal matching property and iteratively achieving optimality. Each iteration of IMF decreases the reverse KL-divergence $\text{KL}(q\|q^{*})$ between the current approximation $q$ and the ground-truth Schrödinger Bridge $q^{*}$ (cf. with IPF). The approach generalizes rectified flows (liu2022flow) to stochastic processes, which are employed (liu2023instaflow; yan2024perflow) in modern foundational generative models such as Stable Diffusion 3 (esser2024scaling). Like IPF, IMF may also accumulate errors. Specifically, it may fail to approximate data distributions due to an imperfect fit at each iteration, causing the marginal matching property to be lost.
@@
In practice, to stabilize IMF training, prevent error accumulation and loss of marginal matching property, practitioners use a heuristic modification of IMF. This is a bidirectional procedure alternating between learning forward and backward processes, either by diffusion-based models in the Diffusion Schrödinger Bridge Matching (DSBM) algorithm (shi2023diffusion) or GANs in Adversarial Schrödinger Bridge Matching (ASBM) algorithm (gushchin2024adversarial). In this work, we investigate the properties of the heuristic modification of the IMF. Our contributions:
@@
These contributions demonstrate that the IPMF procedure has significant potential to unify a range of previously introduced SB methods—including IPF and IMF-based ones—in both discrete (gushchin2024adversarial; de2021diffusion) and continuous time (shi2023diffusion; peluchetti2023diffusion; vargas2021solving) settings, as well as their online versions (de2024schr; peluchetti2024bm; karimi2024sinkhorn). Furthermore, the forward-backward IPMF framework could enable rectified flows to avoid error accumulation, making them even more powerful in generative modeling.
@@ 2 Background
Recall that the SB problem (Schrodinger1931umkehrung), IPF, and IMF admit both discrete– and continuous–time setups leading to the same problem solution. Moreover, the explicit formulas for IPF and IMF in the discrete setting are expressed in terms of probability densities, which helps to convey the main idea of our paper. Thus, for the sake of presentation flow, the main text focuses exclusively on the discrete setup, while Appendix C presents the continuous setup.
@@ 2.1 Schrödinger Bridge (SB) Problem
The SB problem with a Wiener prior in the discrete-time setting (de2021diffusion), given the initial distribution $p_{0}(x_{0})$ and the final distribution $p_{1}(x_{1})$, is stated as
@@
where $\Pi(p_{0},p_{1})\subset\mathcal{P}_{2,ac}(\mathbb{R}^{D\times D})$ is the subset of joint distributions $q(x_{0},x_{1})$ s.t. $q(x_{0})=p_{0}(x_{0})$, $q(x_{1})=p_{1}(x_{1})$. Finally, the static SB objective can be expanded (gushchin2023entropic, Eq. 7)
@@
that is equivalent to the objective of the entropic optimal transport (EOT) problem with the quadratic cost up to an additive constant (cuturi2013sinkhorn; peyre2019computational; leonard2013survey; genevay2019entropy).
@@ 2.2 Iterative Proportional Fitting (IPF)
Early works on SB (vargas2021solving; de2021diffusion; tang2024simplified) propose computational methods based on the IPF procedure (kullback1968probability). The IPF-based algorithm is started by setting the process $q^{0}(x_{0},x_{\text{in}},x_{1})=p_{0}(x_{0})p^{W^{\epsilon}}(x_{\text{in}},x_{1}|x_{0})$. Then, the algorithm alternates between two types of IPF projections, $\text{proj}_{1}$ and $\text{proj}_{0}$, given by (de2021diffusion, Prop. 2):
@@
Thus, $\text{proj}_{1}$ and $\text{proj}_{0}$ replace marginal distributions $q(x_{1})$ and $q(x_{0})$ in $q(x_{0},x_{\text{in}},x_{1})$ by $p_{1}(x_{1})$ and $p_{0}(x_{0})$ respectively. The constructed sequence $\{q^{k}\}$ converges to the solution of the SB problem $q^{*}$ and causes the forward KL-divergence $\text{KL}(q^{*}\|q^{k})$ to decrease monotonically at each iteration. In practice, since the prior process $p^{W^{\epsilon}}$ is used only for initialization, the imperfect fit may lead to a deviation from the SB solution at some iteration. This problem is called “prior forgetting” and was discussed in (vargas2024transport, Appx. E.3). The authors of vargas2021solving consider a continuous analog of the IPF procedure using inversions of diffusion processes (see Appx. C.2).
@@ 2.3 Iterative Markovian Fitting (IMF)
The Iterative Markovian Fitting (IMF) procedure (peluchetti2023diffusion; shi2023diffusion; gushchin2024adversarial) emerged as a strong competitor to the IPF procedure. In contrast to IPF, IMF does not suffer from the “prior forgetting”. The procedure is initialized with any $q^{0}\in\Pi_{N}(p_{0},p_{1})$. Then it alternates between reciprocal projection $\text{proj}_{\mathcal{R}}$ and Markovian projection $\text{proj}_{\mathcal{M}}$:
@@
The reciprocal projection $\text{proj}_{\mathcal{R}}$ creates a new (in general, non-Markovian) process combining the distribution $q(x_{0},x_{1})$ and $p^{W^{\epsilon}}(x_{\text{in}}|x_{0},x_{1})$. The latter is called the discrete Brownian Bridge. The Markovian projection $\text{proj}_{\mathcal{M}}$ uses the set of transitional densities $\{q(x_{t_{n}}|x_{t_{n-1}})\}$ or $\{q(x_{t_{n}}|x_{t_{n+1}})\}$ to create a new Markovian process starting from $q(x_{0})$ or $q(x_{1})$ respectively. Markovian projection keeps the marginal distributions at each timestep, but, in general, changes the joint distributions between them. The sequence $\{q^{k}\}$ converges to the SB $q^{*}$ and causes the reverse KL-divergence $\text{KL}(q^{k}\|q^{*})$ to decrease monotonically at each iteration (cf. with IPF). The authors of shi2023diffusion; peluchetti2023diffusion consider a continuous-time version of the IMF.
@@ 2.4 Heuristic bidirectional modification of IMF
The result of the Markovian projection (7) admits both forward and backward representation. To learn the corresponding transitional densities, one uses neural networks $\{q_{\theta}(x_{t_{n}}|x_{t_{n-1}})\}$ (forward parametrization) or $\{q_{\phi}(x_{t_{n}}|x_{t_{n+1}})\}$ (backward parametrization). The starting distributions are as follows: $q_{\theta}(x_{0})=p_{0}(x_{0})$ for the forward parametrization and $q_{\phi}(x_{1})=p_{1}(x_{1})$ for the backward parametrization. In practice, the alternation between representations of Markovian processes is used in both implementations of continuous-time IMF by DSBM algorithm (shi2023diffusion, Alg. 1) based on diffusion models and discrete-time IMF by ASBM algorithm (gushchin2024adversarial, Alg. 1) based on the GANs. This bidirectional procedure can be described as follows:
@@ 3.1 Bidirectional IMF is IPMF
The heuristic bidirectional IMF alternates between two IMF projections ($\text{proj}_{\mathcal{M}}(\text{proj}_{\mathcal{R}}(\cdot))$) during which the process “became more optimal” (step towards optimality property) and two IPF projections (proj0 and proj1) during which the marginal fitting improves (step towards marginal matching property). We refer to this procedure as Iterative Proportional Markovian Fitting (IPMF). An IPMF step consists of two IMF projections and two IPF projections. We hypothesize that IPMF converges from any initial process $q^{0}(x_{0},x_{\text{in}},x_{1})$, unlike IPF and IMF, which require a specific form of the starting process. We emphasize that IPMF reduces to IMF when the initial coupling has the correct marginals $p_{0}$ and $p_{1}$ and has Brownian Bridge between the marginals. Similarly, if the initial coupling is Markovian, is in the reciprocal class, and has the correct initial marginal $p_{0}$ or $p_{1}$, then IPMF reduces to IPF. Fig. 1 visualizes these cases, clarifying the role of the initial coupling and the iterative steps. A similar analysis for continuous-time IPMF is provided in Appx. C.3.
@@ 3.2 Theoretical Convergence Analysis In Various Cases
Problem (11) is the OT problem with the transport cost $c_{A}(x_{0},x_{1}):=-x_{1}^{\top}Ax_{0}$ and entropy regularization (with weight $1$) (cuturi2013sinkhorn; genevay2019entropy). In other words, for any $2D$ Gaussian distribution $q$, there exists a matrix $A(q)\in\mathbb{R}^{D\times D}$ that defines the cost function for which $q$ solves the EOT problem. We name $A(q)$ the optimality matrix. If $q$ is such that $A(q)=\epsilon^{-1}I_{D}$, then the corresponding transport cost is $c_{A}(x_{0},x_{1})=-\epsilon^{-1}\cdot\langle x_{1},x_{0}\rangle$ which is equivalent to $\epsilon^{-1}\cdot\|x_{1}-x_{0}\|^{2}/2$. Consequently, $q$ is the static SB (2) between its $q_{0}(x_{0})$ and $q_{1}(x_{1})$ for the prior $W^{\epsilon}$, recall (3).
@@
An IPF step does not change the “copula”, i.e., the information about the joint distribution that is invariant w.r.t. changes in marginals $p_{0}$ and $p_{1}$. In the Gaussian case can be represented by the optimality matrix $A_{k}$ (see Lemma D.3). In contrast, an IMF iteration changes the copula but preserves the marginals. Next, we analyze closed formulas for the IMF step in the Gaussian case (peluchetti2023diffusion; gushchin2024adversarial) and show that the IMF step makes $A_{k}$ closer to $\epsilon^{-1}I_{D}$. Specifically, we verify the contractivity of each step w.r.t. $A_{k}$. ∎
@@
Assume $p_{0}$ and $p_{1}$ have bounded supports. Then $q^{4k}(x_{0},x_{1}\!)\!\stackrel{{\scriptstyle w}}{{\rightarrow}}\!q^{*}(x_{0},x_{1}\!)$}, i.e., the discrete-time IPMF weakly conv. to $q^{*}$.
@@
Related works and our novelty. Our work provides the first theoretical analysis of bidirectional IMF, whereas prior studies analyzed only vanilla IMF. (shi2023diffusion, Theorem 8) and (gushchin2024adversarial, Theorem 3.6) proved sublinear convergence of IMF in reverse KL divergence for continuous and discrete cases. For IPF, sublinear convergence in forward KL divergence under mild assumptions, as well as geometric convergence for Gaussians, are shown in (de2021diffusion, Propositions 4 and 43). Previous results cannot be directly generalized to IPMF. First, IPF and IMF converge in different divergence measures, and a decrease in one does not imply a reduction in the other. Second, IPF updates marginals at each step, so IMF must optimize toward a moving target, whereas pure IMF has a fixed optimum. Unlike (shi2023diffusion), which proves convergence only from the IPF starting coupling, our analysis applies to arbitrary starting couplings. We also view the starting coupling as a tunable hyperparameter and examine its effect in the next section.

### v5

@@ Abstract
The Iterative Markovian Fitting (IMF) procedure, which iteratively projects onto the space of Markov processes and the reciprocal class, successfully solves the Schrödinger Bridge (SB) problem. However, an efficient practical implementation requires a heuristic modification—alternating between fitting forward and backward time diffusion at each iteration. This modification is crucial for stabilizing training and achieving reliable results in applications such as unpaired domain translation. Our work reveals a close connection between the modified version of IMF and the Iterative Proportional Fitting (IPF) procedure—a foundational method for the SB problem, also known as Sinkhorn’s algorithm. Specifically, we demonstrate that the heuristic modification of the IMF effectively integrates both IMF and IPF procedures. We refer to this combined approach as the Iterative Proportional Markovian Fitting (IPMF) procedure. Through theoretical and empirical analysis, we establish the convergence of the IPMF procedure under various settings, contributing to developing a unified framework for solving SB problems. Moreover, from a practical standpoint, the IPMF procedure enables a flexible trade-off between image similarity and generation quality, offering a new mechanism for tailoring models to specific tasks. The code for our method can be found at: https://github.com/gregkseno/ipmf.
@@ 1 Introduction
Diffusion Bridge models inspired by the Schrödinger Bridge (SB) theory, which connects stochastic processes with optimal transport, have recently become powerful approaches in biology (Tong et al., 2024; Bunne et al., 2023), chemistry (Somnath et al., 2023; Igashov et al., ; Kim et al., 2024), image processing (Liu et al., 2023a; Shi et al., 2023; Korotin et al., ; Gushchin et al., 2024a) and speech processing (Chen et al., 2023b). Most of these applications deal with either supervised domain translation, e.g., image super-resolution and inpainting (Liu et al., 2023a) or with unpaired translation, e.g., image style-transfer (Shi et al., 2023) or single-cell data analysis (Tong et al., 2024).
@@
This work specifically focuses on unpaired domain translation (Zhu et al., 2017, Fig. 2). In this setting, given two domains represented solely by unpaired samples, the goal is to transform a sample from the input domain into a sample related to it in the target domain. In this context, researchers usually use SB-based algorithms because they enforce two key properties: the optimality property, ensuring similarity between the input and the translated object, and the marginal matching property, ensuring the translation of the input domain to the target domain. The motivation for relying on such specialized methods, rather than general text-to-data models, is further discussed in Appx. B.
@@
Early works (De Bortoli et al., 2021; Vargas et al., 2021; Chen et al., 2021; Pavon et al., 2021) on using the SB for unpaired domain translation employed the well-celebrated Iterative Proportional Fitting (IPF) procedure (Kullback, 1968), also known as the Sinkhorn algorithm (Cuturi and Doucet, 2014). The IPF procedure is initialized with a simple prior process satisfying the optimality property. It then refines this process iteratively through optimality-preserving transformations until the marginal matching property is achieved. In each iteration, IPF decreases the forward KL-divergence $\text{KL}(q^{*}\|q)$ between the current approximation $q$ and the ground-truth Schrödinger Bridge $q^{*}$. However, in practice, approximation errors may cause IPF to suffer from the “prior forgetting”, where the marginal matching property is achieved but optimality is lost (Vargas et al., 2024; 2021).
@@
The Iterative Markovian Fitting (IMF) procedure (Shi et al., 2023; Peluchetti, 2023a; Gushchin et al., 2024b; Ksenofontov and Korotin, 2025) emerged as a promising competitor to IPF. Contrary to IPF, IMF starts from a stochastic process satisfying the marginal matching property and iteratively achieving optimality. Each iteration of IMF decreases the reverse KL-divergence $\text{KL}(q\|q^{*})$ between the current approximation $q$ and the ground-truth SB $q^{*}$ (cf. with IPF). The approach generalizes rectified flows (Liu et al., 2022) to stochastic processes, which are employed (Liu et al., 2023b; Yan et al., 2024) in modern foundational models such as Stable Diffusion 3 (Esser et al., 2024). Like IPF, IMF may also accumulate errors. Specifically, it may fail to approximate data distributions due to an imperfect fit at each iteration, causing the marginal matching property to be lost.
@@
In practice, to stabilize IMF training, prevent error accumulation and loss of marginal matching property, practitioners use a heuristic modification of IMF. This is a bidirectional procedure alternating between learning forward and backward processes, either by diffusion-based models in the Diffusion Schrödinger Bridge Matching (DSBM) algorithm (Shi et al., 2023) or GANs in Adversarial Schrödinger Bridge Matching (ASBM) algorithm (Gushchin et al., 2024b). In this work, we investigate the properties of the heuristic modification of the IMF. Our contributions:
@@
These contributions demonstrate that the IPMF procedure has significant potential to unify a range of previously introduced SB methods—including IPF and IMF-based ones—in both discrete (Gushchin et al., 2024b; De Bortoli et al., 2021) and continuous time (Shi et al., 2023; Peluchetti, 2023a; Vargas et al., 2021) settings, as well as their online versions (De Bortoli et al., 2024; Peluchetti, 2025; Karimi et al., 2024). Furthermore, the forward-backward IPMF framework could enable rectified flows to avoid error accumulation, making them even more powerful in generative modeling.
@@ 2 Background
Recall that the SB problem (Schrödinger, 1931), IPF, and IMF admit both discrete– and continuous–time setups leading to the same problem solution. Moreover, the explicit formulas for IPF and IMF in the discrete setting are expressed in terms of probability densities, which helps to convey the main idea of our paper. Thus, for the sake of presentation flow, the main text focuses exclusively on the discrete setup, while Appendix C presents the continuous setup.
@@ 2.1 Schrödinger Bridge (SB) Problem
The SB problem with a Wiener prior in the discrete-time setting (De Bortoli et al., 2021), given the initial distribution $p_{0}(x_{0})$ and the final distribution $p_{1}(x_{1})$, is stated as
@@
where $\Pi(p_{0},p_{1})\subset\mathcal{P}_{2,ac}(\mathbb{R}^{D\times D})$ is the subset of joint distributions $q(x_{0},x_{1})$ s.t. $q(x_{0})=p_{0}(x_{0})$, $q(x_{1})=p_{1}(x_{1})$. Finally, the static SB objective can be expanded (Gushchin et al., 2023a, Eq. 7)
@@
that is equivalent to the objective of the entropic optimal transport (EOT) problem with the quadratic cost up to an additive constant (Cuturi, 2013; Peyré et al., 2019; Léonard, 2013; Genevay, 2019).
@@ 2.2 Iterative Proportional Fitting (IPF)
Early works on SB (Vargas et al., 2021; De Bortoli et al., 2021; Tang et al., 2024) propose computational methods based on the IPF procedure (Kullback, 1968). The IPF-based algorithm is started by setting the process $q^{0}(x_{0},x_{\text{in}},x_{1})=p_{0}(x_{0})p^{W^{\epsilon}}(x_{\text{in}},x_{1}|x_{0})$. Then, the algorithm alternates between two types of IPF projections, $\text{proj}_{1}$ and $\text{proj}_{0}$, given by (De Bortoli et al., 2021, Prop. 2):
@@
Thus, $\text{proj}_{1}$ and $\text{proj}_{0}$ replace marginal distributions $q(x_{1})$ and $q(x_{0})$ in $q(x_{0},x_{\text{in}},x_{1})$ by $p_{1}(x_{1})$ and $p_{0}(x_{0})$ respectively. The constructed sequence $\{q^{k}\}$ converges to the solution of the SB problem $q^{*}$ and causes the forward KL-divergence $\text{KL}(q^{*}\|q^{k})$ to decrease monotonically at each iteration. In practice, since the prior process $p^{W^{\epsilon}}$ is used only for initialization, the imperfect fit may lead to a deviation from the SB solution at some iteration. This problem is called “prior forgetting” and was discussed in (Vargas et al., 2024, Appx. E.3). The authors of Vargas et al. (2021) consider a continuous analog of the IPF procedure using inversions of diffusion processes (see Appx. C.2).
@@ 2.3 Iterative Markovian Fitting (IMF)
The Iterative Markovian Fitting (IMF) procedure (Peluchetti, 2023a; Shi et al., 2023; Gushchin et al., 2024b) emerged as a strong competitor to the IPF procedure. In contrast to IPF, IMF does not suffer from the “prior forgetting”. The procedure is initialized with any $q^{0}\in\Pi_{N}(p_{0},p_{1})$. Then it alternates between reciprocal projection $\text{proj}_{\mathcal{R}}$ and Markovian projection $\text{proj}_{\mathcal{M}}$:
@@
The reciprocal projection $\text{proj}_{\mathcal{R}}$ creates a new (in general, non-Markovian) process combining the distribution $q(x_{0},x_{1})$ and $p^{W^{\epsilon}}(x_{\text{in}}|x_{0},x_{1})$. The latter is called the discrete Brownian Bridge. The Markovian projection $\text{proj}_{\mathcal{M}}$ uses the set of transitional densities $\{q(x_{t_{n}}|x_{t_{n-1}})\}$ or $\{q(x_{t_{n}}|x_{t_{n+1}})\}$ to create a new Markovian process starting from $q(x_{0})$ or $q(x_{1})$ respectively. Markovian projection keeps the marginal distributions at each timestep, but, in general, changes the joint distributions between them. The sequence $\{q^{k}\}$ converges to the SB $q^{*}$ and causes the reverse KL-divergence $\text{KL}(q^{k}\|q^{*})$ to decrease monotonically at each iteration (cf. with IPF). The authors of Shi et al. (2023); Peluchetti (2023a) consider a continuous-time version of the IMF.
@@ 2.4 Heuristic bidirectional modification of IMF
The result of the Markovian projection (7) admits both forward and backward representation. To learn the corresponding transitional densities, one uses neural networks $\{q_{\theta}(x_{t_{n}}|x_{t_{n-1}})\}$ (forward parametrization) or $\{q_{\phi}(x_{t_{n}}|x_{t_{n+1}})\}$ (backward parametrization). The starting distributions are as follows: $q_{\theta}(x_{0})=p_{0}(x_{0})$ for the forward parametrization and $q_{\phi}(x_{1})=p_{1}(x_{1})$ for the backward parametrization. In practice, the alternation between representations of Markovian processes is used in both implementations of continuous-time IMF by DSBM algorithm (Shi et al., 2023, Alg. 1) based on diffusion models and discrete-time IMF by ASBM algorithm (Gushchin et al., 2024b, Alg. 1) based on the GANs. This bidirectional procedure can be described as follows:
@@ 3.1 Bidirectional IMF is IPMF
The heuristic bidirectional IMF alternates between two IMF projections ($\text{proj}_{\mathcal{M}}(\text{proj}_{\mathcal{R}}(\cdot))$) during which the process “became more optimal” (step towards optimality property) and two IPF projections (proj0 and proj1) during which the marginal fitting improves (step towards marginal matching property). We refer to this procedure as Iterative Proportional Markovian Fitting (IPMF). An IPMF step consists of two IMF projections and two IPF projections. We hypothesize that IPMF converges from any initial process $q^{0}(x_{0},x_{\text{in}},x_{1})$, unlike IPF and IMF, which require a specific form of the starting process. We emphasize that IPMF reduces to IMF when the initial coupling has the correct marginals $p_{0}$ and $p_{1}$ and has Brownian Bridge between the marginals. Similarly, if the initial coupling is Markovian, is in the reciprocal class, and has the correct initial marginal $p_{0}$ or $p_{1}$, then IPMF reduces to IPF. Fig. 1 visualizes these cases, clarifying the role of the initial coupling and the iterative steps. A similar analysis for continuous-time IPMF is provided in Appx. C.3 and the summarization of the conceptual differences between prior bidirectional SB heuristics and the proposed IPMF framework is presented in Table 3.
@@ 3.2 Theoretical Convergence Analysis In Various Cases
Problem (11) is the OT problem with the transport cost $c_{A}(x_{0},x_{1}):=-x_{1}^{\top}Ax_{0}$ and entropy regularization (with weight $1$) (Cuturi, 2013; Genevay, 2019). In other words, for any $2D$ Gaussian distribution $q$, there exists a matrix $A(q)\in\mathbb{R}^{D\times D}$ that defines the cost function for which $q$ solves the EOT problem. We name $A(q)$ the optimality matrix. If $q$ is such that $A(q)=\epsilon^{-1}I_{D}$, then the corresponding transport cost is $c_{A}(x_{0},x_{1})=-\epsilon^{-1}\cdot\langle x_{1},x_{0}\rangle$ which is equivalent to $\epsilon^{-1}\cdot\|x_{1}-x_{0}\|^{2}/2$. Consequently, $q$ is the static SB (2) between its $q_{0}(x_{0})$ and $q_{1}(x_{1})$ for the prior $W^{\epsilon}$, recall (3).
@@
An IPF step does not change the “copula”, i.e., the information about the joint distribution that is invariant w.r.t. changes in marginals $p_{0}$ and $p_{1}$. In the Gaussian case can be represented by the optimality matrix $A_{k}$ (see Lemma D.3). In contrast, an IMF iteration changes the copula but preserves the marginals. Next, we analyze closed formulas for the IMF step in the Gaussian case (Peluchetti, 2023a; Gushchin et al., 2024b) and show that the IMF step makes $A_{k}$ closer to $\epsilon^{-1}I_{D}$. Specifically, we verify the contractivity of each step w.r.t. $A_{k}$. ∎
@@
Assume $p_{0}$ and $p_{1}$ have bounded supports. Then for both discrete-time and continuous-time IPMF it holds $q^{4k}(x_{0},x_{1}\!)\!\stackrel{{\scriptstyle w}}{{\rightarrow}}\!q^{*}(x_{0},x_{1}\!)$}, where $\stackrel{{\scriptstyle w}}{{\rightarrow}}$ denotes weak convergence.
@@
Related works and our novelty. Our work provides the first theoretical analysis of bidirectional IMF, whereas prior studies analyzed only vanilla IMF. (Shi et al., 2023, Theorem 8) and (Gushchin et al., 2024b, Theorem 3.6) proved sublinear convergence of IMF in reverse KL divergence for continuous and discrete cases. For IPF, sublinear convergence in forward KL divergence under mild assumptions, as well as geometric convergence for Gaussians, are shown in (De Bortoli et al., 2021, Propositions 4 and 43). Previous results cannot be directly generalized to IPMF. First, IPF and IMF converge in different divergence measures, and a decrease in one does not imply a reduction in the other. Second, IPF updates marginals at each step, so IMF must optimize toward a moving target, whereas pure IMF has a fixed optimum. Unlike (Shi et al., 2023), which proves convergence only from the IPF starting coupling, our analysis applies to arbitrary starting couplings. We also view the starting coupling as a tunable hyperparameter and examine its effect in the next section.